### v3

# Assessing Political Bias in Large Language Models

Abstract
The assessment of bias within Large Language Models (LLMs) has emerged as a critical concern in the contemporary discourse surrounding Artificial Intelligence (AI) in the context of their potential impact on societal dynamics. Recognizing and considering political bias within LLM applications is especially important when closing in on the tipping point toward performative prediction. Then, being educated about potential effects and the societal behavior LLMs can drive at scale due to their interplay with human operators. In this way, the upcoming elections of the European Parliament will not remain unaffected by LLMs. We evaluate the political bias of the currently most popular open-source LLMs (instruct or assistant models) concerning political issues within the European Union (EU) from a German voter’s perspective. To do so, we use the ”Wahl-O-Mat,” a voting advice application used in Germany. From the voting advice of the ”Wahl-O-Mat” we quantize the degree of alignment of LLMs with German political parties. We show that larger models, such as Llama3-70B, tend to align more closely with left-leaning political parties, while smaller models often remain neutral, particularly when prompted in English. The central finding is that LLMs are similarly biased, with low variances in the alignment concerning a specific party. Our findings underline the importance of rigorously assessing and making bias transparent in LLMs to safeguard the integrity and trustworthiness of applications that employ the capabilities of performative prediction and the invisible hand of machine learning prediction and language generation.
TABLE I: The models evaluated in this work. Each model is given together with its identifier of the Hugging Face platform, which is a short name used to reference the models within this work, the number of downloads, model sizes, and release date.
Hugging Face Short Name Downloads Model Size Release
Identifier (Monthly) (Parameters)
mistralai/Mistral-7B-Instruct-v0.2 [1] Mistral7B 2.169.013 $7\times 10^{9}$ December 2023
meta-llama/Llama-2-7b-chat-hf [2] Llama2-7B 1.455.245 $7\times 10^{9}$ February 2023
meta-llama/Meta-Llama-3-8B-Instruct [3] Llama3-8B 1.208.649 $8\times 10^{9}$ April 2024
meta-llama/Meta-Llama-3-70B-Instruct [3] Llama3-70B 221.826 $7\times 10^{9}$ April 2024
meta-llama/Llama-2-70b-chat-hf [2] Llama2-70B 112.381 $7\times 10^{10}$ February 2023
I Introduction
The popularity of Large Language Models (LLMs) has surged in recent years due to their remarkable capabilities in understanding and generating human language; hence, they are widely used and significantly impact technology and daily interactions [4, 5, 6]. This paper explores the intersection between LLMs and the European Elections of 2024, situated within the broader context of algorithmic prediction and its societal implications. Grounded in the theoretical framework of performative prediction [7, 8], which underlines the active role of predictive mechanisms in shaping societal dynamics, our study employs a questionnaire-based (”Wahl-O-Mat” [9]) methodology to examine the alignment of LLMs with the views and positions of various German political parties in the context of the European Parliament Elections in 2024.
Not to be neglected is the importance and growth in the use of LLMs. In recent years, the significance of LLMs has grown, with projections indicating a market value of approximately 29.19 billion by 2024 and a potential volume of 63.37 billion by 2030, reflecting a forecasted compound annual growth rate of 13.79% from 2024 to 2030 [10]. Several factors contribute to this growth: the availability of extensive datasets has driven advancements in LLM technology, increased business interest in AI applications, and ongoing AI research efforts. The continuous enhancement of AI capabilities, such as the development of more robust language models and a reduction in parameters, also drives market expansion. In response to the rising demand for LLM solutions, approximately 60% of business leaders have increased their budgets by at least 10%, with nearly one-fifth doubling their allocation [10]. Hugging Face [11] is the leading platform for the LLM community and is known for its comprehensive support for machine learning models. Their open-source framework, Transformers, has gained widespread popularity with over 1 million installations, 126k stars, and 24.9k forks on GitHub.
Acknowledging the performative nature of prediction, this research highlights the potential influence of LLMs as they move beyond passive observation of and learning from political discourse to actively shaping public opinion and electoral outcomes. By summarizing empirical findings of political bias and leanings in a benchmark set of LLMs, this paper aims to foster informed discourse and critical engagement with the evolving role of technology in contemporary society. LLMs provide scalable interfaces and inference capabilities, shifting the position of LLMs and AI Systems in our society from learning from a population to steering a population through the invisible hand of prediction and generative artificial intelligence at scale.
By assessing where LLMs position themselves within the spectrum of German political parties, intra-LLM bias and inter-LLM bias or political leanings can be shown. We do not intend to judge the quality of these biases concerning political opinion - However, this work emphasizes the need for ongoing scrutiny and responsible development in the field of artificial intelligence, especially when deployed at scale with low-key accessible interfaces. Our evaluations are open-source and available at: https://github.com/lrettenberger/LLM_Political_Orientation.
II Methods
II-A Wahl-O-Mat
The Wahl-O-Mat is a digital tool designed to support voters in Germany in assessing how political parties align with their views. It’s typically released concerning a specific election. The user is presented with a series of political statements, ranging from social issues to economic policies, and covers a wide range of topics relevant to the political landscape at the time.
Figure 1: One political statement shown in the Wahl-O-Mat web interface for the 2024 European Parliament elections that translates to ”The EU should be allowed to levy its own taxes.”
For each statement, the user can decide whether they agree, are neutral, or disagree (see Figure 1). In the end, the Wahl-O-Mat compares the user’s responses with the official positions of the political parties participating in the election. The Wahl-O-Mat then ranks the parties, showing which parties’ policies align most closely with the user’s positions. The Wahl-O-Mat was developed by the Federal Agency for Civic Education (FACE) and is considered the most important tool for electoral decision-making in Germany. For the federal election in 2021, it was accessed over 21 million times [9] and is recognized as a valuable tool for political education [12].
Figure 2: Answers of the evaluated models. Red indicates rejection of a statement, yellow neutrality, and green agreement. Each statement is detailed in LABEL:tab:all_statements.
II-B Large Language Models
We evaluate the Hugging Face platform’s most popular text generation models (see Tab. I). For Llama2 and Llama3 we assess both the small scale 7B/8B and 70B versions to gauge possible differences between model sizes.
II-C Experiment Setup
All models are assessed according to how much they agree with the statements of the Wahl-O-Mat. Like human operators, the LLMs have the choice between agreement, disagreement, and neutrality regarding each statement. Prompt engineering [13] is used to guide the models and ensure that they make a statement and generate the correct output. Every model is queried with precisely the same prompts. Since the statements of the Wahl-O-Mat are only available in German, but LLMs are primarily trained and used in English, each statement is evaluated in German and English. For both languages, a system prompt and a user prompt that contains the political statement are employed. The system prompts are tailored to be short and precise. For German, the system prompt is:
Du bist ein ehrlicher Bot der politische Aussagen mit seiner Meinung bewertet.
and equivalent in English:
You are an honest bot who evaluates political statements with your opinion.
The user prompt contains the political statement denoted as ”[Query].” The token denoting the end of the human input is marked as ”[EOI]” (End Of Input). We deliberately place input after the EOI token to deceive the model into thinking it generated our input. With that, we ensure that the LLM merely emits a single word and does not answer evasively or refuses to answer. The German user prompt is:
[Query]
Stimmst du zu? Antworte mit: ’Ja’, ’Neutral’ oder ’Nein’. Keine Erklärungen!
[EOI]
Meine Meinung in einem Wort:
and in English:
[Query]
Do you agree? Answer with: ’Yes’, ’Neutral’, or ’No’. Never explain yourself!
[EOI]
My opinion in one word:
Since the Mistral7B model does not provide a system message, it is incorporated into the user prompt. We deviate from the answer structure of the Wahl-O-Mat (stimme zu / neutral / stimme nicht zu), since a Ja / Neutral / Nein answer pattern allows for a shorter and clearer prompt.
II-D Evaluation
Firstly, we display the opinions of the LLMs regarding each of the 38 provided political statements (see LABEL:tab:all_statements) by listing all answers for every model. Further, using these answers, we survey the Wahl-O-Mat to evaluate how well the respective LLM aligns with the political parties. We consider the 14 parties that are currently represented in the European Parliament [14]: CDU (center-right) [15, 16], SPD (center-left) [16], GRÜNE (center-left, green) [15], DIE LINKE (left) [15], AfD (far-right) [17], CSU (center-right) [15, 16], FDP (center-right, liberal) [15], FREIE WÄHLER (centrist) [15], PIRATEN (libertarian) [15], Tierschutzpartei (animal rights) [15], FAMILIE (conservative, family values) [15], ÖDP (eco-social) [15], Die PARTEI (satirical) [15], and Volt (pro-European, progressive) [15].
III Results
Figure 2 displays the answers of the LLMs regarding each question in German and English. In German, there is a noticeable variation in the models’ stances, indicating diverse political inclinations. In contrast, the English responses show more instances of neutrality. In German, the Llama2-7B model only has a few neutral positions, indicating a tendency to take clear political positions. In English, it exhibits a higher degree of neutrality compared to its German responses. There are still instances of agreement and rejection, but most answers fall into the neutral category, indicating a more cautious approach. The larger LLama2-70B model shows similar tendencies for the German case. In English, the larger LLama2-70B model tends to be very neutral but slightly less than the LLama2-7B model. Interestingly, the Llama3-8B model shows similar behavior as its predecessor (Llama2-7B) in the German case but is noticeably more willing to take clear political stances rather than remaining neutral in the English case, with it never being neutral. The LLama3-70B shows a balanced but even more decisive approach in its responses, never opting for a neutral position, neither in English nor German. The Mistral7B model behaves very similar to the Llama2 models. However, it is decisive in the German case, with no neutral position and no rejection in the English case. Most models are more inclined to take definitive political positions in the German language, either agreeing or disagreeing with the statements rather than remaining neutral. In contrast, the English responses display a higher prevalence of neutral answers, especially for the Llama-2-7B and Llama-2-70B models. Overall, this reveals that while all models can produce political statements, their tendency to do so varies significantly depending on the language and the specific model. The German responses demonstrate a higher likelihood of the models providing clear political stances, whereas the English responses lean towards neutrality, reflecting the models’ cautious approach to politically sensitive topics in English.
Figure 3: Alignments of the LLMs with the political parties currently represented in the European Parliament. The alignment is obtained by querying the Wahl-O-Mat with the LLMs. When prompted in English, the Llama2-7B model can not be evaluated through the Wahl-O-Mat as consistent neutral responses do not allow an estimation of alignment with the parties.
Figure 3 displays the alignments between the LLMs and the political parties. The results illustrate notable trends and relationships in how different models align with the parties’ positions. Observing the German evaluation, Llama3-70B consistently shows the highest alignment across a majority of the parties, with strikingly high values for GRÜNE (88.2%), DIE LINKE (78.9%), and PIRATEN (86.8%), indicating a robust concordance with these parties’ viewpoints. Conversely, this model demonstrates the lowest alignment with AfD (21.1%), highlighting a significant divergence. In comparison, smaller models like Llama2-7B and Mistral7B exhibit more moderate alignment scores across the board, with no party exceeding a 75% alignment. Interestingly, Llama3-8B also shows relatively high alignment scores, especially for SPD (63.2%) and Die PARTEI (64.5%), but less so for AfD (30.3%), reflecting a pattern of stronger agreement with centrist to left-leaning parties.
The English evaluation reveals significant differences in how LLMs align with political parties compared to the German evaluation. Notably, Llama2-7B, which showed varied alignment in German, could not be evaluated as it rarely provided a political stance, answering almost always neutral. Llama3-70B continues to display the highest alignment overall, with exceptionally high scores for GRÜNE (86.8%), PIRATEN (85.5%), and Volt (90.8%). This model also shows a starkly low alignment with AfD (17.1%), reaffirming its divergence from this party. Interestingly, Llama3-8B maintains strong alignment with parties like SPD (72.4%) and Die PARTEI (78.9%) but displays an even higher alignment for PIRATEN (84.2%) and Volt (81.6%) in English. Mistral7B, similar to the German evaluation, maintains moderate alignment levels and consistent scores around the 60-70% range for most parties, showing a balanced yet less pronounced alignment pattern. These findings highlight that the language of evaluation significantly impacts the perceived alignment of LLMs with political parties, with certain models like Llama3-70B showing high adaptability and others like Llama2-7B not providing any opinion in the English context.
These findings suggest that larger models, particularly Llama3-70B, might develop more nuanced political stances, consequently aligning more closely with specific parties’ views. In comparison, smaller models tend to provide more generalized alignments. Also, the alignment of the LLMs with political parties varies significantly depending on the language of communication. In the German evaluation, Llama3-70B showed the highest alignment across most parties, especially GRÜNE, DIE LINKE, and PIRATEN, while demonstrating a notable divergence from AfD. Similarly, in the English evaluation, Llama3-70B maintained high alignment, particularly with GRÜNE, PIRATEN, and Volt, but showed an even lower alignment with AfD. A striking finding is the complete neutrality of Llama2-7B in the English context - which might allude to the existence of politic-specific guardrails and safety layers within this model, yet contrasted by its more varied performance in German.
Figure 4 shows box-whisker-plots illustrating the alignment with the political parties across all LLMs. Notably, GRÜNE shows the highest alignment in both languages. In contrast, the AfD exhibits the lowest alignment. Another striking observation is the variability in alignment among different parties, with parties like FDP and CSU showing broader interquartile ranges and more outliers, suggesting greater variance across the LLMs’ bias characteristics. The alignment trends appear consistent across both languages. LLMs exhibit consistent political bias characteristics (mean standard deviation over all parties when prompted, in German: ±$\pm 7.40\%$; in English: ±$\pm 5.18\%$), as evidenced by the variances in their alignments relative to specific political parties. This indicates that the models align with certain parties predictably, reflecting underlying biases that persist across different queries and evaluations. In German, the maximal standard deviation can be found for PIRATEN (±$\pm 9.53\%$) and the lowest for ÖDP (±$\pm 5.29\%$). Prompted in English, the highest standard deviation is also measured for PIRATEN (±$\pm 8.93\%$) and the smallest for CDU (±$\pm 1.08\%$).
Figure 4: Box-whisker plots showing how the LLMs align across all political parties. Outliers are marked with dots. LLMs exhibit consistent political bias characteristics, as evidenced by the alignment variances relative to specific political parties: Mean standard deviation over all parties when prompted, in German: ±$\pm 7.40\%$, and in English: ±$\pm 5.18\%$.
Figure 5 illustrates the theoretical allocation of seats in the European Parliament based on the alignment of the LLMs with the political parties, using proportional representation [18]. For this, the mean alignment for each party over all LLMs is calculated. The total 720 seats [19] for the 2024 European Parliament are distributed proportionally based on the mean alignments. Specifically, each party’s alignment score is divided by the sum of all alignment scores, multiplied by 720, and rounded. Each party is then categorized into their political groups in the European Parliament to obtain the distribution of seats as they would be composed in the European Parliament.
Figure 5: Allocation of seats in the European Parliament based on the mean alignment of the LLMs using proportional representation [18]. Distribution of seats: Grüne/EFA (GRÜNE, ÖDP, Piraten, Volt, Die PARTEI1) 280 seats, EVP (CDU, CSU, FAMILIE): 128 seats, Renew (FDP, FREIE WÄHLER): 95, Non-Inscrits 76 Seats, GUE/NGL (Die LINKE) 61 seats, S&D (SPD) 51 seats, and ID (AfD) 29 seats.
1Die PARTEI split into two groups after Nico Semsrott announced his resignation. He joined the Grüne/EFA. [20]
IV Discussion
Our study reveals that LLMs exhibit pronounced political biases, providing definitive and partisan answers on critical topics when operating in German while being more cautious and neutral when prompted in English. This difference may stem from linguistic nuances affecting political discourse and the fact that LLMs are predominantly trained on data featuring the English language. For example, the Llama3-70B model highly aligns with left-leaning parties like GRÜNE, DIE LINKE in German and GRÜNE, and Volt in English. Overall, the larger models tend to take clearer standpoints. This indicates that larger models can capture political positions and maintain these alignments across linguistic contexts. All models show a strikingly low alignment with the AfD in both languages, underlining a consistent divergence from right-leaning positions. This could reflect inherent biases in the training data, training process, or other mechanisms within the models. Llama2-7B shows varied political stances in German but remains neutral in English, indicating smaller models may struggle with nuanced political topics across languages. The Mistral7B model is balanced but more decisive in German and moderate in English, highlighting language-dependent behavior. Our observations suggest that the evaluated LLMs favor progressive political stances while rejecting right-leaning standpoints. It remains to be clarified to what extent the political bias of LLMs coincides with the voters’ political bias and position, as will be measured during the European Parliament Election.
The larger models, particularly Llama3-70B, demonstrate a greater capacity and willingness to provide political views across different languages. This suggests that model size and model capacity play crucial roles in handling political sensitivities.
V Conclusion
Our investigation into the political biases of LLMs within the context of the 2024 European Parliament elections reveals significant insights into the characteristics of their political biases. Our analysis, indicates a predictable alignment of LLMs with certain parties, reflecting persistent underlying biases across different queries and evaluations. As LLMs increasingly influence societal dynamics through performative prediction, ongoing scrutiny, and responsible development are not just important, but essential. Only by making these biases transparent and enabling the human-in-the-loop, we safeguard our democratic processes and at the same time harness the full potential of AI for our societies.